A gas in a cylinder is expanded from to at a constant pressure of 1 MPa. Determine the work done in this process.
1
Expert's answer
2021-09-24T13:25:30-0400

Assume that "V_1=1\ (m^3)" and "V_2=6\ (m^3)" . So,


"W=p\cdot\Delta V=1 000000\cdot(6-1)=5000000\ (J)=5\ (MJ)"